# TAP4LLM: Table Provider on Sampling, Augmenting, and Packing Semi-structured Data for Large Language Model Reasoning

### 14 Dec 2023 v1

@@
Abstract.
Table reasoning has shown remarkable progress in a wide range of table-based tasks. These challenging tasks require reasoning over both free-form natural language (NL) questions and semi-structured tabular data. However, previous table reasoning solutions suffer from significant performance degradation on “huge” tables. In addition, most existing methods struggle to reason over complex questions since they lack essential information or they are scattered in different places. To alleviate these challenges, we exploit a table provider, namely TAP4LLM, on versatile sampling, augmentation, and packing methods to achieve effective semi-structured data reasoning using large language models (LLMs), which 1) decompose raw tables into sub-tables with specific rows/columns based on the rules or semantic similarity; 2) augment table information by extracting semantic and statistical metadata from raw tables while retrieving relevant knowledge from trustworthy knowledge sources (e.g., Wolfram Alpha, Wikipedia); 3) pack sampled tables with augmented knowledge into sequence prompts for LLMs reasoning while balancing the token allocation trade-off. We show that TAP4LLM allows for different components as plug-ins, enhancing LLMs’ understanding of structured data in diverse tabular tasks.
xxx Reference Format:
xxx, 14(1): XXX-XXX, 2024.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@xx.org. Copyright is held by the owner/author(s). Publication rights licensed to the xx Endowment.
Proceedings of the xx Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
xxx Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://anonymous.4open.science/r/TableProvider-4CC3.
Figure 1. Demonstration of the Three TAP4LLM Modules.
@@ Abstract.
1. Introduction
The extensive and complex characteristics of data are commonly represented in the format of structured data. It makes the data more manageable and facilitates data analysis and processing by machines. Table is one of those fundamental and widely used semi-structured data types in relational databases, spreadsheet applications, and programming languages that handle data for various domains, including financial analysis (Zhang et al., 2020; Li et al., 2022), risk management (Babaev et al., 2019), healthcare analytics (Vamathevan et al., 2019), etc. Reasoning over tabular data is a fundamental aspect of natural language understanding (NLU) and information retrieval (IR), and has several downstream tasks, such as Table-based Question Answering (TQA) (Chen et al., 2020b; Iyyer et al., 2017; Ye et al., 2023; Cheng et al., 2023), Table-based Fact Verification (TFV) (Chen et al., 2020a; Xie et al., 2022; Günther et al., 2021), Table-to-Text (Wang et al., 2021a), Text-to-SQL (Yu et al., 2018), Column Type & Relation Classification (Iida et al., 2021; Deng et al., 2020),etc.
In the field of natural and programming language understanding and generation, LLMs (large language models) have demonstrated impressive abilities as few-shot reasoners with prompt engineering and in-context learning (Kojima et al., 2022a, b). Their capacity to generate human-like text and human-level code (Cheng et al., 2023; Ye et al., 2023; Gemmell and Dalton, 2023) has opened up new possibilities for reasoning over tabular data (Chen, 2022; Gong et al., 2020). However, it is non-trivial to directly utilize vanilla LLMs such as GPT (Liu et al., 2023) to consume tables as their prompt inputs.
@@ 1. Introduction
First, which part of a table should be kept in the prompt? The full content of a table could be very long and noisy to be included in the prompt. Most LLMs have a limited input context window size (e.g., 4k tokens) in which an overlong table cannot fit. For long tables that satisfy the length constraint, it can still lead to unnecessary computations (of LLM on long prompt) and quality regressions (generation interfered by noisy input) when placing irrelevant table content (w.r.t. the task or query) in the prompt. To address the challenge, some table sampling methods were proposed in ad-hoc ways. For example, truncating the input tables to contain only the first 22 rows and 8 columns (Chen, 2022), or filtering relevant rows based on $n$-gram overlap between them and the utterance (Yin et al., 2020). To answer the question of which part to keep, we need a more systematic study of different grounding and sampling algorithms.
@@
Second, what additional/external knowledge could help LLMs better understand a table? The raw content of a table may contain ambiguous information (e.g., abbreviations, domain-specific terms, column type,etc.) that requires further interpretation and clarification. As a result, direct reasoning on the raw tables may lead to misinterpretation and bias in the LLMs’ outputs. To address this, some augmentation techniques were proposed to incorporate structured knowledge (Sui et al., 2023; Xie et al., 2022), common sense knowledge (Bian et al., 2023; Ogundare and Araya, 2023; Shen et al., 2023; Guo et al., 2023), and analytical knowledge (Jena et al., 2022; He et al., 2023) into pre-training and inference processes. For example, (Jena et al., 2022) transforms existing tabular data to create diverse natural language inference instances for better zero-shot performance. (He et al., 2023) infers implicit metadata behind raw table content through field distribution and knowledge graph information. However, the techniques were proposed independently and there lack of comprehensive study that compares them and tries to combine them to provide useful and diverse knowledge and thoughts for LLMs.
@@
Third, how do we encode the table into a prompt? While sampling and grounding compress the table content, augmentation expands the prompt by adding more information. With a given token budget, one needs to find the balance to allocate available tokens between table content and augmented knowledge. Furthermore, the serialization format of the table also plays a critical role. It not only influences how well an LLM understands the table input (as discussed in our previous study  (Sui et al., 2023)), but also determines the string length of the serialized table and the augmented information. For example, as discussed in (Sui et al., 2023), table formats such as HTML (Aghajanyan et al., 2021) or XML are better understood by GPT models, but they also lead to increased token consumption. To pack a table into the prompt, these problems should be addressed with trade-offs.
@@
In this paper, we propose TAP4LLM (table provider for large language models) as a versatile pre-processing toolbox to generate table prompts in LLM reasoning. TAP4LLM addresses the above challenges with three corresponding modules (i) Table Sampling: Selecting a sub-table $T^{\prime}$ from the raw table $T$ based on the rules or semantics of the query or utterance $Q$; (ii) Table Augmentation: Enhancing $T^{\prime}$ by integrating relevant external knowledge, metadata, and attributes based on the raw Table $T$; and (iii) Table Packing: Packing the sampled table $T^{\prime}$ with the augmented knowledge into a sequence with a specified serialization format for LLMs while balancing the token allocation trade-off. In each module, we collect and design many different methods for usage in various scenarios (e.g., speed over accuracy). Across six distinct datasets, our findings demonstrate that TAP4LLM significantly improves accuracy by achieving an average enhancement of 6.02% through the table sampling module, 3.29% through the table augmentation module, and 1.38% through the table packing module. Collectively, TAP4LLM elevates accuracy by an average of 7.93% when compared to the direct input of raw tables into LLMs.. Our exploration has led us to conclude that TAP4LLM enhances the interaction between LLMs and tabular data by employing effective pre-processing.
In addition, through multiple empirical studies, we develop a comprehensive guideline for the effective utilization of TAP4LLM. Our investigation reveals that the optimal performance of sampling is achieved through the query-based method. We then conducted evaluations on augmentation methods and proposed the most suitable combination of sampling and augmentation for each dataset. Ablation studies are also conducted on token allocation to identify trade-offs between model performance and resource usage.
@@
  • We proposed a unified pre-processor TAP4LLM to improve the effectiveness of LLMs in tabular reasoning tasks. Our framework includes three modules: table sampling, table augmentation, and table packing.
  • We conducted a comprehensive survey to evaluate the effectiveness of TAP4LLM. Through various sampling, augmentation, and packing methods, TAP4LLM achieves an average of improved performance by 7.93%. We also assessed the effectiveness of different configurations inside TAP4LLM, including embedding types and statistical features.
  • We formulated a comprehensive usage guideline for our framework TAP4LLM. First, we identify the optimal combination of methods within TAP4LLM for different usage scenarios. Second, we provide a trade-off map between performance metrics and token allocation.
2. Preliminaries
Figure 2. TAP4LLM Framework for Tabular Data. Noted that the “table sync” refers to the application (such as Excel Copilot) keeps its table data in sync with the table manager. The table manager acts as an intermediary, managing the data that is either stored locally in a cache or accessed through a database connection. This syncing process is crucial for “interactive table reasoning” and for maintaining data integrity. The implication of this syncing process is further discussed in §5.
Table Reasoning Tasks. Each instance in table-based reasoning consists of a table $T$, a natural language question $Q$, and an answer $A$. Specifically, table $T$ is defined as $T=\left\{v_{i,j}\mid i\leq\operatorname{R}_{T},j\leq\operatorname{C}_{T}\right\}$, containing $R_{T}$ rows and $C_{T}$ columns. The content of the cell in the $i$-th row and $j$-th column is represented by $v{i,j}$. A question $Q$ is a sequence of $n$ tokens: $Q=\{q_{1},q_{2},q_{3},\cdots,q_{n}\}$. In this paper, our primary focus is on two distinct table-based reasoning tasks, table-based fact verification (TFV) and table-based question answering (TQA). In TFV, the answer $A$ is a boolean value in $\{0,1\}$, indicating the veracity of the input statement (where $1$ means the statement is entailed by the given table, and 00 means the statement is refuted by the given table). In TQA, the answer is a sequence of natural language tokens represented as $A=\{a_{1},a_{2},a_{3},\cdots,a_{n}\}$ corresponding to the posed question. For our experiments, all tables first undergo table sampling and table augmentation by our proposed method and then are serialized into a sequence by table packing and serialization. Note that various serialization methods have already been explored in our previous work (Sui et al., 2023). Detailed implementation specifics are provided in Section §3.3.
Challenges with Long Sequences in LLMs. Most LLMs struggle with long sequences due to the quadratic complexity of self-attention mechanisms (Vaswani et al., 2017; Tay et al., 2022). It’s important to note that the maximum sequence length of commonly used LLMs, such as gpt-3.5-turbo-16k (zot, [n.d.]), is limited to 16k tokens. This limit becomes especially problematic when dealing with structured data, which typically consists of various components and causes significant memory and computational challenges. Solutions like Tapex (Liu et al., 2022) and Tapas (Herzig et al., 2020) utilize a simple truncation approach based on the table’s sequence length. However, these methods exhibit the shortcomings of excluding important data and disrupting the whole table structure. In our work, we introduce specific constraints for the LLMs’ call requests. For example, to avoid structural disruptions from truncation, we use a “token allocation” strategy to estimate the token capacity and initiate the sampling functions (explained in Section §3.1) when the table’s token count surpasses a certain threshold.
Neuro-Symbolic Functions & Metadata Analysis. Neuro-symbolic method is a specialized approach in Deep Learning (DL) that combines neural and symbolic methodologies (Hitzler et al., 2022). In this work, we use Metadata Analysis (He et al., 2023), a neuro-symbolic method based on a transformer architecture (Vaswani et al., 2017), to generate metadata essential for table analysis. Such metadata serves as an explicit representation of formally organized background knowledge. An example of this is the binary metadata classification for column types, known as dimension and measure [^1] A measure is a table field consisting of numerical data that you can calculate, such as ’Price’ and ’Discount’. Conversely, a dimension is a field that offers categorical information used for filtering, grouping, and labeling, such as the ’Product Name’ and ’Category’. It’s important to correctly classify fields as measures or dimensions because they dictate the operations you can perform on the data, affecting the accuracy and relevance of the analysis.. Furthermore, various LLM-centric toolsets have been developed to facilitate: (i) integration of LLMs with diverse data sources, and (ii) dynamic interactions of LLMs with their surroundings. We select some of symbolic functions from the toolsets as our agents for table augmentation, including Langchain (LangChain, 2022), Semantic Kernel (matthewbolanos, 2023), etc.
3. TAP4LLM: Table Provider for LLMs
@@
3.1. Table Sampling
One of the primary challenges for reasoning over tabular data with LLMs is (1) token efficiency. Due to the inherent structure of table(s), even a modest-sized table can quickly consume a significant portion of the available token budget of the LLMs. While study(Sui et al., 2023) shows that representing tables in formats like HTML (Aghajanyan et al., 2021) or XML can improve the performance of GPT models, it also leads to increased token usage. Furthermore, (2) noisy information becomes an issue in large tables, especially those with over 30 rows, and becomes worse when the row count exceeds 60 (Ye et al., 2023; Cheng et al., 2023). Such noise can obscure the primary intent of a query, leading to potentially skewed or irrelevant model outputs. Even with advancements like GPT-4222gpt-4-32k: maintains the same capabilities as standard GPT-4 mode but has a 4x context length, equating to 32,768 tokens. More details can be found at https://platform.openai.com/docs/models/gpt-4., which can handle up to 32k tokens, the challenges caused by noisy information remain (Gemmell and Dalton, 2023). In this context, although more and more LLMs have been developed to handle longer sequences, the complexity of reasoning over tabular data highlights the need for sampling techniques. These approaches focus on the most relevant part of a table, providing a practical solution to bridge the gap between the nature of tables and the context boundaries of LLMs.
@@ 3.1. Table Sampling
In table sampling, a subset of top-ranked rows and columns is selected to form the sub-table. Specifically, given an original table $T$ with a distinct set of rows $R_{T}$, columns $C_{T}$, and a query $q$, the goal of table sampling is to produce a sub-table $T^{\prime}=T_{r,c}$, where $r\in\mathcal{P}(R_{T})$, $c\in\mathcal{P}(C_{T})$. Here $\mathcal{P}(X)$ denotes the power set of $X$, representing all possible subsets of $X$. The table sampling formulation can be represented as follows:
(1) $T^{\prime}=T_{r,c}=select(T,rank(f(T,q)))$
The $f(T,q)$ function represents each sampling method. For example, the query-based sampling (discussed in detail below) calculates the similarity score as $f$ between the query $q$ and each row/column from $T$. The $rank()$ function sorts the rows and columns of $T$ based on sampling methods $f$ and outputs a ranked list. The $select()$ function then chooses the top-k rows and top-l columns from the ranked list to form the sub-table $T_{r,c}$. Note that, in the context of column-based sampling, the term “grounding” is frequently used. Grounding becomes essential when dealing with ambiguous queries. For instance, the term “date” could refer to calendar data, a romantic outing, or the fruit. If a table includes columns like “release date”, “date of birth”, or “date ate”, grounding helps determine which column(s) the query refers to (See the improvement of using column grounding in Table 4). Specifically, we propose two distinct variants for table sampling as follows:
3.1.1. Rule-based Sampling

### 17 Feb 2024 v2

@@
Abstract
Table-based reasoning has shown remarkable progress in combining deep models with discrete reasoning, which requires reasoning over both free-form natural language (NL) questions and semi-structured tabular data. However, previous table reasoning solutions only consider small-sized tables and exhibit limitations in handling larger tables. In addition, most existing methods struggle to reason over complex questions since they lack essential information or they are scattered in different places. To alleviate these challenges, we propose TAP4LLM as a versatile pre-processing toolbox to generate table prompts through (1) table sampling, (2) table augmentation, and (3) table packing while balancing the token allocation trade-off. In each module, we collect and design several common methods for usage in various scenarios (e.g., speed over accuracy). We also provide a comprehensive evaluation on performance of each components inside TAP4LLM and show that our method improves LLMs’ reasoning capabilities in various tabular tasks and enhances the interaction between LLMs and tabular data by employing effective pre-processing. The code of this paper will be released at https://anonymous.4open.science/r/TableProvider-4CC3.
[…]
TAP4LLM: Table Provider on Sampling, Augmenting, and Packing Semi-structured Data for Large Language Model Reasoning
Yuan Sui1thanks: The contributions by Yuan Sui, Jiaru Zou, and Xinyi He have been conducted and completed during their internships at Microsoft Research Asia, Beijing, China., Jiaru Zou2***, Mengyu Zhou3thanks: Corresponding author., Xinyi He4***, Lun Du3, Shi Han3, Dongmei Zhang3 1 National University of Singapore 2 University of Illinois Urbana-Champaign 3 Microsoft 4 Xi’an Jiaotong University yuan.sui@u.nus.edu, jiaruz2@illinois.edu, hxyhxy@stu.xjtu.edu.cn {mezho, lun.du, shihan, dongmeiz}@microsoft.com
1 Introduction
The extensive and complex characteristics of data are commonly represented in the format of structured data. Table is one of those fundamental and widely used semi-structured data types in relational databases, spreadsheet applications, and programming languages that handle data for various domains, including financial analysis Zhang et al. (2020); Li et al. (2022), risk management Babaev et al. (2019), healthcare analytics Vamathevan et al. (2019), etc. Reasoning over tabular data is a fundamental aspect of natural language understanding (NLU) and information retrieval (IR), and has several downstream tasks, such as Table-based Question Answering (TQA) Chen et al. (2020b); Iyyer et al. (2017); Ye et al. (2023a); Cheng et al. (2023), Table-based Fact Verification (TFV) Chen et al. (2020a); Xie et al. (2022); Günther et al. (2021), Table-to-Text Wang et al. (2021b), Text-to-SQL Yu et al. (2018), Column Type & Relation Classification Iida et al. (2021); Deng et al. (2020),etc.
Figure 1: Demonstration of the Three TAP4LLM Modules.
@@ 1 Introduction
Meanwhile, large language models (LLMs) are advancing in their ability to tackle challenges associated with natural language using in-context learning Cheng et al. (2023); Ye et al. (2023a); Gemmell and Dalton (2023), but the degree to which they understand tables and how to leverage LLMs to work with table-based data remain an open question Chen (2022); Gong et al. (2020). Our research aims to explore this question and improve how LLMs use and work with table-based data.
@@
First, which part of a table should be kept in the prompt? The full content of a table could be very long and noisy to be included in the prompt. Most LLMs have a limited input context window size (e.g., 4k~16k tokens) in which an overlong table cannot fit. For long tables that satisfy the length constraint, it can still lead to unnecessary computations (of LLM on long prompt) and quality regressions (generation interfered by noisy input) when placing irrelevant table content (w.r.t. the task or query) in the prompt. To address the challenge, some sampling methods were proposed in ad-hoc ways. For example, truncating the input tables to contain only the first 20+ rows and 8 columns Chen (2022), or filtering relevant rows based on $n$-gram overlap between them and the utterance Yin et al. (2020). To answer the question of which part to keep, we need a more systematic study of different grounding and sampling algorithms.
@@
Second, what additional/external knowledge could help LLMs better understand a table? The raw content of a table may contain ambiguous information (e.g., abbreviations, domain-specific terms, column type,etc.) that requires further interpretation and clarification. As a result, direct reasoning on the raw tables may lead to misinterpretation and bias in the LLMs’ outputs. To address this, some augmentation techniques were proposed to incorporate structured knowledge Sui et al. (2023); Xie et al. (2022), common sense knowledge Bian et al. (2023); Ogundare and Araya (2023); Shen et al. (2023); Guo et al. (2023), and analytical knowledge Jena et al. (2022); He et al. (2023) into pre-training and inference processes. For example, Jena et al. (2022) transforms existing tabular data to create diverse NL inference instances for better zero-shot performance. He et al. (2023) infers implicit metadata behind raw table content through field distribution and knowledge graph information. However, the techniques were proposed independently and there lack a comprehensive study that compares them and tries to combine them to provide useful and diverse knowledge and thoughts for LLMs.
@@
Third, how do we encode the table into a prompt? While sampling and grounding compress the table content, augmentation expands the prompt by adding more information. With a given token budget, one needs to find the balance to allocate available tokens between table content and augmented knowledge. Furthermore, the serialization format of the table also plays a critical role. It not only influences how well an LLM understands the table input (as discussed in our previous study  Sui et al. (2023)), but also determines the string length of the serialized table and the augmented information. For example, as discussed in Sui et al. (2023), table formats such as HTML Aghajanyan et al. (2021) or XML are better understood by GPT models, but they also lead to increased token consumption. To pack a table into the prompt, these problems should be addressed with trade-offs.
@@
In this paper, we propose TAP4LLM (table provider for large language models) as a versatile pre-processing toolbox to generate table prompts in LLM reasoning. TAP4LLM addresses the above challenges with three corresponding modules (i) Table Sampling: Selecting a sub-table $T^{\prime}$ from the raw table $T$ based on the rules or semantics of the query or utterance $Q$; (ii) Table Augmentation: Enhancing $T^{\prime}$ by integrating relevant external knowledge, metadata, and attributes based on the raw Table $T$; and (iii) Table Packing: Packing the sampled table $T^{\prime}$ with the augmented knowledge into a sequence with a specified serialization format for LLMs while balancing the token allocation trade-off. In each module, we collect and design several common methods for usage in various scenarios (e.g., speed over accuracy). Across six distinct datasets, our findings demonstrate that TAP4LLM significantly improves accuracy by achieving an average enhancement of 6.02% through the table sampling module, 3.29% through the table augmentation module, and 1.38% through the table packing module. Collectively, TAP4LLM elevates accuracy by an average of 7.93% when compared to the direct input of raw tables into LLMs. Our exploration has led us to conclude that TAP4LLM enhances the interaction between LLMs and tabular data by employing effective pre-processing.
Figure 2: TAP4LLM Framework for Tabular Data. Noted that the “table sync” refers to the application (such as Excel Copilot) keeps its table data in sync with the table manager. The table manager acts as an intermediary, managing the data that is either stored locally in a cache or accessed through a database connection. This syncing process is crucial for “interactive table reasoning” and for maintaining data integrity. The implication of this syncing process is further discussed in §F.2.
@@
  • We proposed a unified pre-processor to improve the effectiveness of LLMs in tabular reasoning tasks and enhance the interaction between LLMs and tabular data.
  • We conducted a comprehensive evaluation of each components and showed that TAP4LLM achieves an average of improved performance by 7.93%.
  • We formulated a complete usage guideline for our framework TAP4LLM. For different real world scenarios, We identify corresponding optimal combination of methods within each module. we also provide a trade-off map between performance metrics and token allocation for reference.
2 TAP4LLM: Table Provider for LLMs
@@
2.1 Table Sampling
@@
In table sampling, a subset of top-ranked rows and columns is selected to form the sub-table. Specifically, given an original table $T$ with a distinct set of rows $R_{T}$, columns $C_{T}$, and a query $q$, the goal of table sampling is to produce a sub-table $T^{\prime}=T_{r,c}$, where $r\in\mathcal{P}(R_{T})$, $c\in\mathcal{P}(C_{T})$. Here $\mathcal{P}(X)$ denotes the power set of $X$, representing all possible subsets of $X$. The process can be formulated as $T^{\prime}=T_{r,c}=\text{select}(T,\text{rank}(f(T,q)))$. The $f(T,q)$ function represents each sampling method. For example, the query-based sampling (discussed in details below) calculates the similarity score as $f$ between the query $q$ and each row / column from $T$. The $rank()$ function sorts the rows and columns of $T$ based on sampling methods $f$ and outputs a ranked list. The $select()$ function then chooses the top-k rows and top-l columns from the ranked list to form the sub-table $T_{r,c}$. Specifically, we classify multiple variants for table sampling as three following categories:
2.1.1 Rule-based Sampling